Suppose you are taking a multiple choice exam with 65 questions, each with 4 possible answers. If you were to guess at random for each question, how many questions would you expect to get right?

Your answer does not need to be an integer.
Answer #1

probability of getting a correct answer for a random question by guessing = p = 1/4 = 0.25

no. of questions = n = 65

Considering the situation to be a Bernoulli process, we get

the expected no. of winnings or right answered questions = n * p = 65 * 0.25 = 16.25

i.e. I am expected to get 16.25 questions with right answers out of 65 questions.